OpenStudy (anonymous):

A culture initially contains 200 bacteria. If the number of bacteria doubles every 2 hours, how many bacteria will be in the culture at the end of 12 hours? a. 2,600 c. 36,300 b. 12,800 d. 3,400

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

can my medals double every two hours?

OpenStudy (anonymous):

b. 12,800 you multiply 200 by 2 six times because it only doubles every 2 hours.

OpenStudy (anonymous):

first lets just do it \[200\\ 400\\ 800\\ 1600\\ 3200\\ 6400\]

OpenStudy (anonymous):

So 12/2=6. Then 200*2*2*2*2*2*2=12800

OpenStudy (anonymous):

oh i get it now

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have a similar question if any1 interested

OpenStudy (anonymous):

what @MLeighW said more succinctly written as \[200\times 2^6\]

OpenStudy (anonymous):

A culture initially contains 200 bacteria. If the number of bacteria doubles every 2 hours, how many bacteria will be in the culture at the end of 12 hours? a. 2,600 c. 36,300 b. 12,800 d. 3,400

OpenStudy (anonymous):

wow looks pretty similar

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Heavy rain caused a river to rise. The river rose three inches the first day, and each additional day it rose twice as much as the previous day. How much did the river rise in five days? a. 102 in c. 93 in b. 145 in d. 15 in

OpenStudy (anonymous):

srry yeah lol

OpenStudy (anonymous):

\[3\times 2^5\] if i am reading it correctly

OpenStudy (anonymous):

it isnt a option

OpenStudy (anonymous):

yeah i see that

OpenStudy (anonymous):

what then????????? HELP some1

OpenStudy (anonymous):

oh i see i think you have to add them

OpenStudy (anonymous):

\[3+6+12+24+48\]

OpenStudy (anonymous):

is 93 an option?

OpenStudy (anonymous):

oh yes thanks yeah